# Critical Masses and Numerical Computation of Massive Scalar Quasinormal Modes in Schwarzschild Black Holes

### v1

@@ Abstract
We present a comprehensive analysis of the quasinormal modes (QNMs) of a massive scalar field in Schwarzschild spacetime using two complementary numerical techniques: the Hill-determinant method and Leaver’s continued-fraction method. Our study systematically compares the performance, convergence, and consistency of the two approaches across a wide range of field masses and angular momenta. We identify three critical mass thresholds, $_{\rm lim}, mm$$_{\rm max}$, and $m_{zd}$, which govern qualitative changes in the QNM spectrum. In particular, long-lived modes emerge at $m_{zd}$, where the imaginary part of the frequency vanishes and the mode becomes essentially non-decaying. This phenomenon is robust across multipoles and may have important implications for the phenomenology of massive fields around black holes. Our results provide a detailed numerical characterization of massive scalar QNMs and highlight the complementary strengths of the Hill-determinant and continued-fraction methods, paving the way for future studies of rotating or charged black holes and quasi-bound states.
@@ I INTRODUCTION
Black holes (BHs) are ideal laboratories for probing the nature of gravity in its most extreme regime. Their response to perturbations is characterized by quasinormal modes (QNMs), damped oscillations whose frequencies are determined exclusively by the background geometry and the nature of the perturbing field. These oscillations dominate the ringdown stage of astrophysical processes, such as binary mergers, and have become a central element in the interpretation of gravitational wave signals[1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
@@
Perturbations of black holes and the associated quasinormal modes (QNMs) have been extensively investigated. Seminal analyses of Schwarzschild stability [11, 12] were soon extended to Reissner–Nordström [13, 14] and Kerr black holes [15], and later to a variety of compact objects within and beyond General Relativity [16, 17, 18, 19, 20, 21, 22]. Beyond their role in gravitational wave astronomy, QNMs are also relevant for astrophysical processes such as the dynamics of accretion disks, active galactic nuclei, and gamma-ray bursts. Following Unruh’s proposal of black hole evaporation analogues in laboratory systems [23], analogue black holes have become a thriving area of research [24, 25, 26, 27, 28].
@@
Accurate computation of quasinormal modes (QNMs) requires dedicated numerical techniques, since the perturbation equations reduce to wave-like forms with effective potentials that lack closed-form solutions. Several approaches have been developed, including direct numerical integration [29], WKB-based methods [30, 31, 32, 33, 34], the asymptotic iteration method [35], and spectral collocation schemes [36]. Among these, two recurrence-based techniques stand out: Leaver’s continued-fraction method [37] and the Hill-determinant method [38]. The continued-fraction approach is widely regarded as the standard, particularly following Nollert’s convergence improvement [39] and subsequent refinements [40]. The Hill-determinant method, though less common, has recently been shown to achieve comparable accuracy with modern convergence accelerators [41].
@@
In this work, we revisit the QNMs of a massive scalar field in Schwarzschild spacetime with two complementary goals. First, we provide a systematic comparison between the Hill-determinant method [38, 42] and Leaver’s continued-fraction method [13, 37] in the massive case, assessing their agreement, stability, and numerical efficiency. Second, we investigate in detail the impact of the scalar field mass on the QNM spectrum. Our analysis reveals the existence of three distinct mass thresholds: $_{\rm lim}, mm$$_{\rm max}$, and $m_{zd}$, which govern qualitative changes in the spectral properties. In particular, we show that at $m_{\rm max}$ the quasinormal spectrum undergoes a qualitative change: while the classical picture suggests the absence of physical modes for $m>m_{\rm max}$, both methods consistently predict the emergence of long-lived modes at $m_{zd}$. These modes, with diverging damping times, persist across multipoles and may have important implications for the phenomenology of massive fields around black holes.
@@ II MASSIVE SCALAR FIELD IN SCHWARZSCHILD SPACETIME
is the effective potential, with l=0,1,2,3l=0,1,2,3… parameterizing the field angular harmonic index. From here on we use units where $2M=1$. It is important to note that, unlike the usual approach, we are not using tortoise coordinates. As a result, Eq. (5) does not take the standard Schrödinger-like form.
@@
There are two special classes of solutions of physical interest: Quasinormal modes (QNM), characterized by $A\left(\omega\right)=0$, corresponding to purely outgoing waves at spatial infinity and Quasibound states (QBS) which are spatially localized and decay exponentially away from the black hole. In both cases, the imposition of appropriate boundary conditions at the horizon and at infinity leads to a discrete spectrum of allowed complex frequencies.
To better understand the differences between quasinormal modes (QNMs) and quasi-bound states (QBSs), and to properly characterize these solutions, it is essential to analyze their behavior in the asymptotic regions. In this context, techniques for solving Eq. (5) reveal that the frequencies $\omega$ generally take complex values of the form
@@
Unlike the massless case, Eq. (5) does not admit only wave-like solutions [^1] By wave-like we mean functions of the form $f(t\pm r/v)$, with $v$ the phase velocity.. For real $\omega^{2}$ and $m^{2}$, propagation requires $\omega^{2}>m^{2}$. In the present setting, however, $\omega^{2}$ is complex, and a more careful analysis is needed. To this end, we define:
@@
  • If $q_{I}>0$, we have a temporal decay $e^{-\omega_{I}t}$ and a spatial growth $e^{q_{I}r}$. This combination describes a damped perturbation that propagates to infinity — a QNM. Physically, this implies a non-zero energy flux at large distances.
@@
An important limiting case is the massless limit. From the definitions of $\rho$ and $\nu$, this corresponds to $m\rightarrow 0$ and $\nu\rightarrow\rho$. Taking the massless limit of Eq. (16), we then obtain:
@@
This yields a massless five-term recurrence relation, which does not allow for a direct comparison with the three-term relation obtained in Ref. [37]. Nevertheless, in Sec. VI we show that both formulations reproduce the same quasinormal mode spectrum.
@@ III THE HILL DETERMINANT METHOD
The Hill determinant method, originally developed in the context of periodic differential equations in mathematical physics, has found significant application in the computation of quasinormal modes (QNMs) of black holes. The method’s adaptation to black hole perturbation theory emerged in the late 1980s, notably with the work of Majumdar and Panchapakesan [38], who demonstrated its effectiveness for determining the complex QNM frequencies of Schwarzschild black holes. More recently, the method has been further refined and extended, incorporating convergence acceleration techniques such as the Wynn algorithm and Borel summation, which have enabled high-precision calculations for higher overtones and for spacetimes in higher dimensions [41, 27, 42, 6]. In particular, Ref. [41] have demonstrated that the Hill determinant approach, combined with double convergence acceleration and Leaver-Nollert-Zhidenko tail approximations, achieves exceptional accuracy and stability, even for modes with small real parts.
@@
These determinants define polynomial equations in $\rho$, which can be systematically analyzed in increasing matrix order. However, imposing the additional condition that the waves be purely outgoing at spatial infinity, along with the requirement of stability, restricts the allowed frequencies $\omega$ to those lying in the complex plane with negative imaginary parts. These correspond precisely to the quasinormal modes. An important point of the Hill determinant method, in contrast to the continued fraction approach Ref. [37], is that it remains applicable to recurrence relations involving more than three terms. Consequently, due to the sparsity of the Hill matrices and the representation of the determinant in the form of Eq. (24), there is no practical need to reduce the original five-term recurrence relation to a tridiagonal form via Gaussian elimination.
@@ IV The Leaver continued fraction method
with the event horizon and spatial infinity now mapped to $r^{*}\to$-$\infty$ and $r^{*}\to+\infty$, respectively. An appropriate ansatz in this case is [43]
@@
Once the three-term recurrence relation is obtained, the standard Leaver’s method can be applied to determine the QNMs as the roots of the algebraic equation [37, 13]
@@
To improve the convergence of the continued fraction on the right-hand side of Eq. (32), we employ the technique developed by Nollert in Ref.[39]. He demonstrated that the convergence of the continued-fraction method is enhanced if the sum is initiated with an appropriate estimate for the rest of the continued fraction, $R_{N}$, defined by
@@ V.1 Roots and behavior near horzion
The Eq. (41) is a cubic equation, and in general, it presents three possible scenarios for its roots, depending on the value of the discriminant $\mathcal{D}$. These scenarios are:
@@
Therefore, we observe that for $r>1$, the function $V(r)$ must either have two real roots greater than $1$ (i.e., it crosses the $x$-axis twice for $r>1$ or no real roots at all (i.e., it never crosses the $x$-axis for $r>1$. Combining this observation with the result above, we conclude that there are no roots for $r>1$, and thus the potential is strictly positive in the region between the horizon and infinity.
@@ V.2 Critical masses
The next step is to study the critical points of the potential. This is important because these points are associated with the critical masses, which in turn produce qualitative changes in the behavior of the quasinormal mode spectrum. By differentiating Eq. (6) by $r$ and setting the result to zero, we obtain:
@@
whose Vieta’s formulas give:
@@
The expression above depends on $m$ and $L$ and can be positive, negative, or zero. In fact, if we fix $L$ and gradually increase $m^{2}$, we see that $\mathcal{D}(m,L)$ transitions from negative (small $m^{2}$) to positive (large $m^{2}$). It is thus useful to solve $\mathcal{D}(m,L)=0$. The critical values $m^{2}_{\text{max}}$ that mark this transition are:
@@
  • For small $m^{2}$, the discriminant $\mathcal{D}(m,L)<0$, and the cubic has two positive roots and one negative root. From Eq. (50), one of the positive roots $r_{3}$ lies deep in the radial domain. The other root $r_{2}$, by continuity from the massless case, lies outside the horizon. Recalling that
@@
    Solving for $m^{2}$, we obtain the critical value:
@@
The limit mass $m_{lim}$ given by Eq. (V.2) has a form similar to that found in Ref. [44]. In that work, the corresponding value is referred to as the maximum mass, defined as the largest mass for which the effective potential still exhibits a peak, and beyond this value, the peak disappears. According to the analysis of QNMs in Ref. [30], quasinormal modes can be interpreted as waves trapped by this peak. Thus, once the peak vanishes, the potential can no longer confine waves, and QNMs cease to exist. In contrast, in the present work, we introduce a different threshold, $m_{\text{lim}}$, defined as the mass at which the potential peak equals its asymptotic value. Thus, for $m>m_{\text{lim}}$, the potential still retains a peak, and we still have QMNs.
@@
Figure 1: The effective potential $V$ as a function of $r$ for $l=0$ (left panel) and $l=1$ (right panel), for differents values of mass $m$.
Fig. 1 shows the behavior of the effective potential for $l=0$ and $l=1$, respectively. As discussed earlier in this section, increasing the mass $m$ causes the potential peak to approach the asymptotic value. At $m=m_{\text{lim}}$, the global maximum becomes clearly a local maximum and transitions into an inflection point when the mass reaches $m=m_{\text{max}}$. For $m>m_{\text{max}}$, the potential peak disappears entirely.
@@ VI.1 Massless limit
As the first part of our analysis, we examine the massless limit and compute the corresponding quasinormal modes (QNMs). This is achieved by solving the five-term recurrence relation given in Eq.(15), with the coefficients defined in Eq.(21), using the method outlined in Sec. III. We compare our results with those obtained via Leaver’s continued fraction method, as presented in Sec. IV. The computed values are summarized in the Table 2.
@@
Table 2 compares quasinormal mode frequencies computed using the Hill determinant method with those obtained via Leaver’s method. The two approaches exhibit excellent agreement, with relative errors consistently small and many cases showing differences within numerical precision. This strong correspondence validates the accuracy of the Hill determinant method for calculating quasinormal modes in Schwarzschild spacetime and confirms its ability to reproduce the well-established spectrum from Leaver’s continued fraction approach. Furthermore, despite our formulation involving a five-term recurrence relation, both methods yield the same quasinormal mode spectrum consistent with the findings of Ref. [31].
Another relevant aspect to highlight is the difference in the distribution of quasinormal modes between the massive and massless cases. Looking at the discussion made in Sec.II, we see that in the massless limit ($m=0$), only quasinormal modes (QNMs) are present. These modes exhibit a characteristic spectral symmetry, as shown in Fig. 2 (left panel). This symmetry was discussed previously in Ref. [45]. In contrast, when a small mass is introduced, this symmetry is broken and quasi-bound states emerge, as illustrated in the right panel of Fig. 2.
Figure 2: Quasi-frequency spectra for the fundamental mode with $l=1$. Left panel: Massless case ($m=0$), showing symmetric quasinormal modes (QNMs) with respect to $\text{Re}(\omega)$. Right panel: Massive case ($m=10^{-5}$), where the symmetry is broken.

### v2

@@ Abstract
We present a comprehensive analysis of the quasinormal modes (QNMs) of a massive scalar field in Schwarzschild spacetime using two complementary numerical techniques: the Hill-determinant method and Leaver’s continued-fraction method. Our study systematically compares the performance, convergence, and consistency of the two approaches across a wide range of field masses and angular momenta. We identify three critical mass thresholds, $_{\rm lim}, mm$$_{\rm max}$, and $m_{zd}$, which govern qualitative changes in the QNM spectrum. In particular, long-lived modes emerge at $m_{zd}$, where the imaginary part of the frequency vanishes and the mode becomes essentially nondecaying. This phenomenon is robust across multipoles and may have important implications for the phenomenology of massive fields around black holes. Our results provide a detailed numerical characterization of massive scalar QNMs and highlight the complementary strengths of the Hill-determinant and continued-fraction methods, paving the way for future studies of rotating or charged black holes and quasibound states.
@@ I INTRODUCTION
Black holes (BHs) are ideal laboratories for probing the nature of gravity in its most extreme regime. Their response to perturbations is characterized by QNMs, damped oscillations whose frequencies are determined exclusively by the background geometry and the nature of the perturbing field. These oscillations dominate the ringdown stage of astrophysical processes, such as binary mergers, and have become a central element in the interpretation of gravitational wave signals[Vishveshwara:1970zz, chandrasekhar1998mathematical, Kokkotas:1999bd, Nollert:1999ji, Berti:2009kk, Konoplya:2011qq, LIGOScientific:2016aoc, LIGOScientific:2017vwq, EventHorizonTelescope:2019dse, EventHorizonTelescope:2022wkp].
@@
Perturbations of black holes and the associated QNMs have been extensively investigated. Seminal analyses of Schwarzschild stability [Regge:1957td, Zerilli:1970wzz] were soon extended to Reissner-Nordström [Leaver:1990zz, Kokkotas:1988fm] and Kerr black holes [Teukolsky:1972my], and later to a variety of compact objects within and beyond general relativity [Correa:2024xki, Molina:2010fb, Blazquez-Salcedo:2016enn, Macedo:2018txb, Berti:2018vdi, Antoniou:2024jku, Zinhailo:2024jzt]. Beyond their role in gravitational wave astronomy, QNMs are also relevant for astrophysical processes such as the dynamics of accretion disks, active galactic nuclei, and gamma-ray bursts. Following Unruh’s proposal of black hole evaporation analogs in laboratory systems [Unruh:1980cg], analog black holes have become a thriving area of research [Dolan:2011ti, Torres:2020tzs, Barcelo:2005fc, Matyjasek:2024uwo, dePaula:2025fqt].
@@
Accurate computation of QNMs requires dedicated numerical techniques, since the perturbation equations reduce to wavelike forms with effective potentials that lack closed-form solutions. Several approaches have been developed, including direct numerical integration [Chandrasekhar:1975zza], WKB-based methods [1985ApJ...291L..33S, Iyer:1986np, Konoplya:2003ii, Matyjasek:2017psv, Galtsov:1991nwq], the asymptotic iteration method [Cho:2009cj], and spectral collocation schemes [Jansen:2017oag]. Among these, two recurrence-based techniques stand out: Leaver’s continued-fraction method [Leaver:1985ax] and the Hill-determinant method [Majumdar:1989tzg]. The continued-fraction approach is widely regarded as the standard, particularly following Nollert’s convergence improvement [Nollert:1993zz] and subsequent refinements [Zhidenko:2006rs]. The Hill-determinant method, though less common, has recently been shown to achieve comparable accuracy with modern convergence accelerators [Benda:2025tni].
@@
In this work, we revisit the QNMs of a massive scalar field in Schwarzschild spacetime with two complementary goals. First, we provide a systematic comparison between the Hill-determinant method [Majumdar:1989tzg, Matyjasek:2021xfg] and Leaver’s continued-fraction method [Leaver:1990zz, Leaver:1985ax] in the massive case, assessing their agreement, stability, and numerical efficiency. Second, we investigate in detail the impact of the scalar field mass on the QNM spectrum. Our analysis reveals the existence of three distinct mass thresholds: $_{\rm lim}, mm$$_{\rm max}$, and $m_{zd}$, which govern qualitative changes in the spectral properties. In particular, we show that at $m_{\rm max}$ the quasinormal spectrum undergoes a qualitative change: while the classical picture suggests the absence of physical modes for $m>m_{\rm max}$, both methods consistently predict the emergence of long-lived modes at $m_{zd}$. These modes, with diverging damping times, persist across multipoles and may have important implications for the phenomenology of massive fields around black holes.
@@ II MASSIVE SCALAR FIELD IN SCHWARZSCHILD SPACETIME
is the effective potential, with l=0,1,2,3l=0,1,2,3… parametrizing the field angular harmonic index. From here on we use units where $2M=1$. It is important to note that, unlike the usual approach, we are not using tortoise coordinates. As a result, Eq. (5) does not take the standard Schrödinger-like form.
@@
There are two special classes of solutions of physical interest: QNM, characterized by $A\left(\omega\right)=0$, corresponding to purely outgoing waves at spatial infinity and quasibound states (QBS) which are spatially localized and decay exponentially away from the black hole. In both cases, the imposition of appropriate boundary conditions at the horizon and at infinity leads to a discrete spectrum of allowed complex frequencies.
To better understand the differences between QNMs and QBSs, and to properly characterize these solutions, it is essential to analyze their behavior in the asymptotic regions. In this context, techniques for solving Eq. (5) reveal that the frequencies $\omega$ generally take complex values of the form
@@
Unlike the massless case, Eq. (5) does not admit only wavelike solutions [^1] By wavelike we mean functions of the form $f(t\pm r/v)$, with $v$ the phase velocity.. For real $\omega^{2}$ and $m^{2}$, propagation requires $\omega^{2}>m^{2}$. In the present setting, however, $\omega^{2}$ is complex, and a more careful analysis is needed. To this end, we define:
@@
  • If $q_{I}>0$, we have a temporal decay $e^{-\omega_{I}t}$ and a spatial growth $e^{q_{I}r}$. This combination describes a damped perturbation that propagates to infinity — a QNM. Physically, this implies a nonzero energy flux at large distances.
@@
An important limiting case is the massless limit. From the definitions of $\rho$ and $\nu$, this corresponds to $m\rightarrow 0$ and $\nu\rightarrow\rho$. Taking the massless limit of Eq. (16), we then obtain
@@
This yields a massless five-term recurrence relation, which does not allow for a direct comparison with the three-term relation obtained in Ref. [Leaver:1985ax]. Nevertheless, in Sec. VI we show that both formulations reproduce the same quasinormal mode spectrum.
@@ III THE HILL DETERMINANT METHOD
The Hill determinant method, originally developed in the context of periodic differential equations in mathematical physics, has found significant application in the computation of QNMs of black holes. The method’s adaptation to black hole perturbation theory emerged in the late 1980s, notably with the work of Majumdar and Panchapakesan [Majumdar:1989tzg], who demonstrated its effectiveness for determining the complex QNM frequencies of Schwarzschild black holes. More recently, the method has been further refined and extended, incorporating convergence acceleration techniques such as the Wynn algorithm and Borel summation, which have enabled high-precision calculations for higher overtones and for spacetimes in higher dimensions [Benda:2025tni, Matyjasek:2024uwo, Matyjasek:2021xfg, Konoplya:2011qq]. In particular, Ref. [Benda:2025tni] have demonstrated that the Hill determinant approach, combined with double convergence acceleration and Leaver-Nollert-Zhidenko tail approximations, achieves exceptional accuracy and stability, even for modes with small real parts.
@@
These determinants define polynomial equations in $\rho$, which can be systematically analyzed in increasing matrix order. However, imposing the additional condition that the waves be purely outgoing at spatial infinity, along with the requirement of stability, restricts the allowed frequencies $\omega$ to those lying in the complex plane with negative imaginary parts. These correspond precisely to the quasinormal modes. An important point of the Hill determinant method, in contrast to the continued fraction approach Ref. [Leaver:1985ax], is that it remains applicable to recurrence relations involving more than three terms. Consequently, due to the sparsity of the Hill matrices and the representation of the determinant in the form of Eq. (24), there is no practical need to reduce the original five-term recurrence relation to a tridiagonal form via Gaussian elimination.
@@ IV The Leaver continued fraction method
with the event horizon and spatial infinity now mapped to $r^{*}\to$-$\infty$ and $r^{*}\to+\infty$, respectively. An appropriate ansatz in this case is [Konoplya:2004wg]
@@
Once the three-term recurrence relation is obtained, the standard Leaver’s method can be applied to determine the QNMs as the roots of the algebraic equation [Leaver:1985ax, Leaver:1990zz]
@@
To improve the convergence of the continued fraction on the right-hand side of Eq. (32), we employ the technique developed by Nollert in Ref.[Nollert:1993zz]. He demonstrated that the convergence of the continued-fraction method is enhanced if the sum is initiated with an appropriate estimate for the rest of the continued fraction, $R_{N}$, defined by
@@ V.1 Roots and behavior near horzion
Equation (41) is a cubic equation, and in general, it presents three possible scenarios for its roots, depending on the value of the discriminant $\mathcal{D}$. These scenarios are
@@
Therefore, we observe that for $r>1$, the function $V(r)$ must either have two real roots greater than $1$ (i.e., it crosses the $x$-axis twice for $r>1$) or no real roots at all (i.e., it never crosses the $x$-axis for $r>1$). Combining this observation with the result above, we conclude that there are no roots for $r>1$, and thus the potential is strictly positive in the region between the horizon and infinity.
@@ V.2 Critical masses
The next step is to study the critical points of the potential. This is important because these points are associated with the critical masses, which in turn produce qualitative changes in the behavior of the quasinormal mode spectrum. By differentiating Eq. (6) by $r$ and setting the result to zero, we obtain
@@
whose Vieta’s formulas give
@@
The expression above depends on $m$ and $L$ and can be positive, negative, or zero. In fact, if we fix $L$ and gradually increase $m^{2}$, we see that $\mathcal{D}(m,L)$ transitions from negative (small $m^{2}$) to positive (large $m^{2}$). It is thus useful to solve $\mathcal{D}(m,L)=0$. The critical values $m^{2}_{\text{max}}$ that mark this transition are
@@
  • For small $m^{2}$, the discriminant $\mathcal{D}(m,L)<0$, and the cubic has two positive roots and one negative root. From Eq. (50), one of the positive roots $r_{3}$ lies deep in the radial domain. The other root $r_{2}$, by continuity from the massless case, lies outside the horizon, and must be negative. Recalling that
@@
    Solving for $m^{2}$, we obtain the critical value
@@
The limit mass $m_{lim}$ given by Eq. (V.2) has a form similar to that found in Ref. [Simone:1991wn]. In that work, the corresponding value is referred to as the maximum mass, defined as the largest mass for which the effective potential still exhibits a peak, and beyond this value, the peak disappears. According to the analysis of QNMs in Ref. [1985ApJ...291L..33S], quasinormal modes can be interpreted as waves trapped by this peak. Thus, once the peak vanishes, the potential can no longer confine waves, and QNMs cease to exist. In contrast, in the present work, we introduce a different threshold, $m_{\text{lim}}$, defined as the mass at which the potential peak equals its asymptotic value. Thus, for $m>m_{\text{lim}}$, the potential still retains a peak, and we still have QMNs.
@@
Figure 1: The effective potential $V$ as a function of $r$ for $l=0$ (left panel) and $l=1$ (right panel), for different values of mass $m$.
Figure 1 shows the behavior of the effective potential for $l=0$ and $l=1$, respectively. As discussed earlier in this section, increasing the mass $m$ causes the potential peak to approach the asymptotic value. At $m=m_{\text{lim}}$, the global maximum becomes clearly a local maximum and transitions into an inflection point when the mass reaches $m=m_{\text{max}}$. For $m>m_{\text{max}}$, the potential peak disappears entirely.
@@ VI.1 Massless limit
As the first part of our analysis, we examine the massless limit and compute the corresponding QNMs. This is achieved by solving the five-term recurrence relation given in Eq.(15), with the coefficients defined in Eq.(21), using the method outlined in Sec. III. We compare our results with those obtained via Leaver’s continued fraction method, as presented in Sec. IV. The computed values are summarized in the Table 2.
@@
Table 2 compares quasinormal mode frequencies computed using the Hill determinant method with those obtained via Leaver’s method. The two approaches exhibit excellent agreement, with relative errors consistently small and many cases showing differences within numerical precision. This strong correspondence validates the accuracy of the Hill determinant method for calculating quasinormal modes in Schwarzschild spacetime and confirms its ability to reproduce the well-established spectrum from Leaver’s continued fraction approach. Furthermore, despite our formulation involving a five-term recurrence relation, both methods yield the same quasinormal mode spectrum consistent with the findings of Ref. [Iyer:1986np].
Another relevant aspect to highlight is the difference in the distribution of quasinormal modes between the massive and massless cases. Looking at the discussion made in Sec.II, we see that in the massless limit ($m=0$), only QNMs are present. These modes exhibit a characteristic spectral symmetry, as shown in Figure 2 (left panel). This symmetry was discussed previously in Ref. [Iyer:1986nq]. In contrast, when a small mass is introduced, this symmetry is broken and quasibound states emerge, as illustrated in the right panel of Figure 2.
Figure 2: Quasifrequency spectra for the fundamental mode with $l=1$. Left panel: massless case ($m=0$), showing symmetric QNMs with respect to $\text{Re}(\omega)$. Right panel: massive case ($m=10^{-5}$), where the symmetry is broken.